An object is embedded in a glass block (n=1.5) 100 cm away from the glass-air interface. The object is observed through the glass by a person standing to the right. The glass-air interface is spherical with radius 20cm with the center of curvature located in air.

An object is embedded in a glass block (n=1.5) 100
Answer #1

Since the object is behind the glass, the distance is negative and we say its virtual.

Using:

\frac{n_1}{o}+\frac{n_2}{i}=\frac{n_1-n_2}{R}

\frac{1.5}{-100[cm]}+\frac{1}{i}=\frac{1.5-1}{-20[cm]}

\frac{1}{i}=-\frac{1.5}{-100[cm]}+\frac{1.5-1}{-20[cm]}

i=-100[cm]

since is positive we say its real and in the right side (front of the glass) so is virtual.

lateral magnification is:

M=-\frac{i}{o}=-\frac{-100}{-100}=-1

If the lateral magnification is negative, the image is inverted.

The refracting surface is diverging.